OpenStudy (anonymous):

(1.) are these functions invertible? if no, say why not, if yes, find the inverse functions. (A.) f(x)=x^3 (B.)h(x)=ln x (C.) t(x)=2x (D.) g(t)=cos t (0degrees

OpenStudy (akshay_budhkar):

A fx=x^3 y=x^3 so the inverse is y=x^1/3 so it is invertible

OpenStudy (akshay_budhkar):

joe

OpenStudy (anonymous):

yes?

OpenStudy (akshay_budhkar):

please help him further..:P

OpenStudy (anonymous):

lol <.<

OpenStudy (anonymous):

h(x) = ln(x) is invertible if you restrict the domain.

OpenStudy (akshay_budhkar):

B h(x)=lnx y=lnx therefore inverse is y=e^x true only in its domain

OpenStudy (anonymous):

t(x) = 2x is invertible no matter what.

OpenStudy (anonymous):

and then cos(t), with 0<t<180 is also invertible.

OpenStudy (akshay_budhkar):

ya the inverse of t(x) is y=x/2

OpenStudy (akshay_budhkar):

i am not sure of it joe??? of cost?

OpenStudy (akshay_budhkar):

procipal domain of cos x is?

OpenStudy (akshay_budhkar):

*principal

OpenStudy (anonymous):

im claim that cos(t) is invertible from 0<t<180 comes from the fact that every value of t will produce a different value of cos(t) in that domain. Its one-to-one (injective).

OpenStudy (anonymous):

^ i do recall htta from precal

OpenStudy (anonymous):

*that

OpenStudy (anonymous):

please help me.

OpenStudy (akshay_budhkar):

yes right joe then it is invertible..

OpenStudy (anonymous):

what is the inverse of g(t) ?

OpenStudy (anonymous):

if the domain had been something like 90<t<270, i would say no , its not invertible

OpenStudy (akshay_budhkar):

the inverse is y=cos^-1(x)

OpenStudy (anonymous):

thank you

OpenStudy (akshay_budhkar):

welcome:)

OpenStudy (akshay_budhkar):

joe say welcome!! n give me a medal

OpenStudy (akshay_budhkar):

lol

OpenStudy (akshay_budhkar):

you slept joe?

OpenStudy (anonymous):

im falling asleep >.<

OpenStudy (akshay_budhkar):

lol i can sense it here!

OpenStudy (anonymous):

just 3 more medals then i lvl up >.<

OpenStudy (anonymous):

just 3 moer questions then im done with my math packet :D